OpenStudy (anonymous):

HELP ME CHECK MY WORK y=4cos(x+(pi/3))-2 a. Amplitude- 4 b. period- 2pi c. phase shift- pi/3 to the left d. vertical translation- down two

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

is it all right @sourwing

OpenStudy (anonymous):

yea

OpenStudy (jdoe0001):

I do think though, the phase shift is negative, is all

OpenStudy (jdoe0001):

\(\bf y=Acos(Bx-C)\qquad \textit{phase shift}=\cfrac{C}{B} \\ \quad \\ y=4cos\left(x+\frac{\pi}{3}\right)\implies y=4cos\left(x-\left[-\frac{\pi}{3}\right]\right)\) other than that, is correct :)